JEE AdvancedProbabilityHard
Question
A box contains 11 tickets numbered from 1 to 11. Two tickets are drawn similtaneously at random.
Let E1 denotes the event that the sum of the numbers on the tickets drawn is even and E2 denotes the event that the sum of the numbers on the tickets drawn is odd
Which of the following hold good?
Let E1 denotes the event that the sum of the numbers on the tickets drawn is even and E2 denotes the event that the sum of the numbers on the tickets drawn is odd
Which of the following hold good?
Options
A.P(E1/E2) = P(E2/E1)
B.E1 and E2 are exhaustive
C.P(E2) > P(E1)
D.E1 and E2 are equally likely
Solution
E1 : Both even or both odd
P(E1) =
P(E2)= 1 - P (E1) =
(i) P(E1 / E2) =
= 0
P(E2 / E1) =
= 0
(ii) E1 and E2 exhaustive
(iii) P(E2) > P(E1)
P(E1) =
P(E2)= 1 - P (E1) =
(i) P(E1 / E2) =
P(E2 / E1) =
(ii) E1 and E2 exhaustive
(iii) P(E2) > P(E1)
Create a free account to view solution
View Solution FreeMore Probability Questions
Two real numbers, x & y are selected at random. Given that 0 ≤ x ≤ 1 ; 0 ≤ y ≤ 1. Let A be the e...Five horses are in a race. Mr. A selects two of the horses at random and bets on them. The probability that Mr. A select...If A and B are two independent events then P (AB) equals-...An unbiased coin is tossed n times. Let X denote the number of times head occurs. If P(X = 4), P (X = 5) and P(X = 6) ar...From a lot containing 10 defective and 90 non-defective bulbs, 8 bulbs are selected one by one with replacement. Then th...